How to convert Celsius in Degrees?Please explain with an example.Thank you

Grade:10

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Amit

K=C+273

if temp is 250c

then in kelvin it will be 25+273=298K

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

1degree celcius = 273kelvin

eg1)27oC = (27+273)=300k

eg2)-5oC = (-5+273)=268K